OpenStudy (lgbasallote):

pls give me some integration problems (no drawings involved) to work on :)

OpenStudy (inkyvoyd):

Ok.

OpenStudy (inkyvoyd):

Watch this one.

OpenStudy (anonymous):

integral of (sec^6 x)/(tan^2 x) dx

OpenStudy (anonymous):

That question came off of one of my midterms

OpenStudy (inkyvoyd):

\[\int\limits_{}^{}e ^{-x^2} dx\]

OpenStudy (inkyvoyd):

\[\int\limits_{}^{}\sin(x^2) dx\]

OpenStudy (inkyvoyd):

\[\int\limits_{}^{}\sin(x^3) dx\]

OpenStudy (inkyvoyd):

And finally, the not so meanie one, \[\int\limits_{}^{}\sec ^3 x dx\]

OpenStudy (anonymous):

\[ \huge\int e^{\sin x}\cdot \frac{x\cos^3x-\sin x}{\cos^2x}dx \]

OpenStudy (inkyvoyd):

\[\int\limits_{}^{}\Gamma(x)erf(x)dx\]

OpenStudy (inkyvoyd):

That last one right there, I'm not sure even if it is valid

OpenStudy (anonymous):

\[\mathsf{\text{Prove }\int\limits_0^{2\pi} f(x) \cos x \ge0,\quad \text{given f''(x)} \ge \text{0 and f(x) is continuous on }{[0,2\pi].}}\]

OpenStudy (anonymous):

An interesting (but somewhat hard) is: Find a real number r and a positive number L such that:\[\LARGE\lim_{x \rightarrow \infty} \frac{r^c \int\limits_{0}^{\pi /2} x^r sinx dx}{\int\limits_{0}^{\pi /2}x^r cosxdx } = L\]

OpenStudy (lgbasallote):

@juanjohnguy \(\LARGE \frac{\tan^3 x}{3} + 2\tan x - \cot x\)

OpenStudy (inkyvoyd):

@bmp , can you do the one's I've given?

OpenStudy (anonymous):

Typo: should be a real number c and a positive number L.

OpenStudy (anonymous):

And the limit is r -> infinity

OpenStudy (inkyvoyd):

@lgbasallote , do sec^3 x, it's a fun integration by parts

OpenStudy (lgbasallote):

lol...how come everytime an asker asks for some problems to work on people give those thingies o.O

OpenStudy (anonymous):

Because they are fun :-)

OpenStudy (inkyvoyd):

You never specified the difficulty.

OpenStudy (lgbasallote):

juan's problem was the only one possible :P

OpenStudy (inkyvoyd):

@bmp , how do you do Integrate[Gamma[x]*Erf[x], x]? Not even mathematica knows

OpenStudy (inkyvoyd):

sec^3 x is very possible. it's integration by parts

OpenStudy (inkyvoyd):

spoiler: u=sec x

OpenStudy (anonymous):

Mine's totally possible. Here's an easier one: \[ \int \frac{dx}{\sqrt{1+e^x}} \]

OpenStudy (inkyvoyd):

Omg, I win. I made the series expansion have the euler-mascheroni constant!

OpenStudy (anonymous):

you almost had it, did you want me to share the answer with you?

OpenStudy (anonymous):

The sin(x^2) is not that hard. Take the imaginary part: \[\ Im( \int_{0}^{\infty} e^{-ix^2}dx ) = \frac{1}{2}\sqrt{\frac{\pi}{2}}\]I think.

OpenStudy (lgbasallote):

lol +C? hahaha @juanjohnguy

OpenStudy (inkyvoyd):

No, it's fresnal s

OpenStudy (inkyvoyd):

try the wolf.

OpenStudy (lgbasallote):

i liked my tutorials better :P

OpenStudy (inkyvoyd):

well, stuf AND fresnel S

OpenStudy (inkyvoyd):

and, isn't that complex analysis or something, @bmp

OpenStudy (anonymous):

\[\mathsf{\text{Evaluate,}\large\int\limits_0^1\left(\sqrt[m]{1-x^n} - \sqrt[n]{1-x^m}\right)\;dx}\]

OpenStudy (anonymous):

I think so, I never took an analysis course. I learned the cos(x^2) and sin(x^2) because they are quite cool :-)

OpenStudy (anonymous):

\[ \int \sec^3x dx=\int(\sec^2x)(\sec x)dx=\tan x \sec x - \int(\tan x)(\sec x \tan x)dx\\ = \tan x \sec x - \int \sec x (\sec^2x-1)dx\\ = \tan x \sec x - \int \sec^3xdx+\int \sec x dx\\ = \frac{1}{2}[\tan x \sec x + \log(\sec x + \tan x)] \]

OpenStudy (anonymous):

lol not quite, the answer that i got was all the same except the beginning is \[(1/2)\tan ^{3}x\]

OpenStudy (lgbasallote):

this thread is laggy :/

OpenStudy (inkyvoyd):

How bout sin(x^3)? why does it have Gamma[] when I mathematica it?

OpenStudy (anonymous):

I almost considered doing a joke relating the laggy thread with the recent change of picture by @lgbasallote but disconsider. @inkyvoyd You got me. I will try to find something :-)

OpenStudy (inkyvoyd):

Alrighty :P

OpenStudy (anonymous):

I can give you another one that was on my old midterm if you like

OpenStudy (lgbasallote):

sure juan..you're the only one who gives possible questions:P haha

OpenStudy (inkyvoyd):

I gave possible questions! hey!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\mathsf{\int \frac{\sec^2x}{\left(\sec x +\tan x\right)^{9/2}}dx}\]

OpenStudy (inkyvoyd):

Wait no. I only gave *a*possible question

OpenStudy (lgbasallote):

only sin^3 x :P

OpenStudy (anonymous):

\[\huge \int x^2 e^xdx\\ \huge\int x^3 e^{x^2} dx\\ \huge\int x^2\sin x dx \] Those are all possible

OpenStudy (anonymous):

A interesting one then \[ \int \sqrt{\tan{x}}dx \]

OpenStudy (inkyvoyd):

Secant ^3 x :P

OpenStudy (inkyvoyd):

I like Integrate[Gamma[x]*Erf[x], x] though

OpenStudy (lgbasallote):

lol50 replies srsly :P @bmp lana gave me that before already haha couldnt solve it

OpenStudy (anonymous):

\[\int\limits_{?}^{?} dx/\sqrt{4x+x^2}\]

OpenStudy (inkyvoyd):

@bmp , is that by parts or u-sub? I could try it out

OpenStudy (anonymous):

@lgbasallote Did you try u = tan(x)? :-)

OpenStudy (anonymous):

the question marks don't mean anything

OpenStudy (anonymous):

\[\mathsf{\large\int\limits_0^{\frac{\pi}2}\left(\sqrt{\ tan x} + \sqrt{\cot x}\right)\;dx}\]

OpenStudy (inkyvoyd):

Wth is PolyGamma?

OpenStudy (anonymous):

I did that once @inkyvoyd . Once. I will never redo it for nothing.

OpenStudy (anonymous):

But it's a u-sub, at least how I solved it.

OpenStudy (lgbasallote):

yes @bmp @juanjohnguy \(\LARGE \frac{1}{2} \ln (\frac{x}{\sqrt{4x + x^2}})\)

OpenStudy (inkyvoyd):

You mean the prob you gave?

OpenStudy (inkyvoyd):

Or wth is polygamma?

OpenStudy (anonymous):

No, the sqrt(tan(x)) one @inkyvoyd . And I think no sane person should work this one out...

OpenStudy (anonymous):

No sorry thats not right. Just a hint is that you have to use trig substitution for that one.

OpenStudy (anonymous):

Feed it to Mathematica and look at the answer. Or use wolfram and check the show steps.

OpenStudy (lgbasallote):

i did..phooey :/

OpenStudy (lgbasallote):

oh i see phooey i removed the sqrt >.<

OpenStudy (anonymous):

lol did you complete the square for the 4x+x^2 part?

OpenStudy (inkyvoyd):

Ok, after I use the bathroom, finish my bio unit test, I will try it.

OpenStudy (inkyvoyd):

sq rt tan x right? Easy to memorize.

OpenStudy (anonymous):

Yeah. But so tricky to solve. :-)

OpenStudy (lgbasallote):

it seems i squared after getting the trig counterparts i forgot it was square rooted..final ansewer \(\Large \ln (\frac{x + 2 + \sqrt{4x + x^2}}{2})\)

OpenStudy (anonymous):

All of it is right according to my answer however the "stuff" under the square root is (x+2)^2 + x^2

OpenStudy (anonymous):

Just cause at the beginning i completed the square for the 4x + x^2

OpenStudy (lgbasallote):

ohh phooey...guess i have a lot to learn haha

OpenStudy (anonymous):

lol but you def almost got it you where extremely close so maybe not too much

OpenStudy (anonymous):

I got the same answer as @lgbasallote :-)

OpenStudy (anonymous):

Here is another \[\int\limits_{}^{}(1+sinx)/(1-sinx) dx\]

OpenStudy (anonymous):

hm....then i guess the integral as 2 answers?

OpenStudy (anonymous):

I mean there are dif ways to do a problem

OpenStudy (anonymous):

I did complete the square first. What did you got afterwards?

OpenStudy (lgbasallote):

i cant solve it T_T i got stuck in cos^2x/(1-sinx)^2

OpenStudy (anonymous):

Maybe u = tan(x/2)? What I got is quite hairy, but it seems solvable.

OpenStudy (anonymous):

after i completed the square i got \[\int\limits_{}^{}dx/\sqrt{(x+2)^2 - 2^2}\] then i made my my x+2= 2 sec\[\theta\] dx=2sec\[\theta \tan \theta\] and the square root of all of the stuff under the square root= 2tan\[\theta\]

OpenStudy (anonymous):

and hold on let me check my paper and see where you are at in relation to mine igbas

OpenStudy (lgbasallote):

i too lagged :/ im closing this thread now

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

perhaps open a new one and it wont be as laggy?